OpenStudy (anonymous):

An airplaine travels against a 52.5 mph wind in 5 hours and makes the return trip with the same wind in 2 hours. find the speed of the plane in still air

OpenStudy (anonymous):

Plane speed w.r.t to wind-a 5(a-52.25)=2(a+52.25)

OpenStudy (anonymous):

They are equal because same distance is travelled

OpenStudy (anonymous):

On left side you are going against wind, right side going with the wind

OpenStudy (anonymous):

what would be the next step

OpenStudy (anonymous):

5(a-52.25)=2(a+52.25) Solve for a

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (anonymous):

distribute 5 into parethesis on left side. Distribute 2 into parenthesis